log 2 8x =18 means that x is equal to ???? please solve this

log2 8x =18 means that x is equal to         
 
???? please solve this

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Hamid
 
log2 8x = 18
 
hence
 
2^18 = 8x
 
2^18 = 2^3 * x
 
x = 2^15
 
Regards
Arun (askIITians forum expert)